Let T:P1→P2T:P1→P2 be a linear transformation defined by

T(a+bx)=3a−2bx+(a+b)x2.T(a+bx)=3a−2bx+(a+b)x2.

(a) Find range(T)range(T) and give a basis for range(T)range(T).

(b) Find ker(T)ker(T) and give a basis for ker(T)ker(T).

(c) By justifying your answer determine whether TT is onto.

(d) By justifying your answer determine whether TT is one-to-one.

(e) Find [T(7+x)]B[T(7+x)]B, where B={−1,−2x,4x2}B={−1,−2x,4x2}.
Answer #1

Range (T) = { T (atba): atba EP, : atba 3 a-aba + (atba ce ? (3+2) a (3+2) a + (-2x+2²) ashari = }(348) a + (-22+x+) b: a be0 kort (at&x) EP, T (arba) -o? { (arba) EP, : (8 + x) a + (-2x + x) (atba) EP, : b(2x-2² 3+22 But then , at bu b(22-0) br + 321-qat 8x2 [1 (10) B = -1, -2*, 4* 1440) T (74x) = 3(7) -2(1) x + (7+1)x² B - (_)(21) + (-2x) 42 (470) [+(7+*)] O).